OpenStudy (anonymous):

can someone help me with my pre-algebra/algebra 1 homework

OpenStudy (bluemiku):

i don't see anything...

OpenStudy (anonymous):

The water tank had a maximum capacity of 84 gallons. If the tank was 2/7 full, how many gallons of water did it have? George cut 2/3 of the pie and put that giant piece on his plate. Then he ate 1/4 of that piece. what fraction of the original pie did George eat? Mr. Drysdale earned $906.25 in interest in one year on money that he deposited in his bank. If the ban paid an interest rate of 6.25%, how much money did Mr. Drysdale deposit? There’s some number that if you subtract from it first, and then multiply that total by 7, the result is 28. Find the number.

OpenStudy (anonymous):

there are some of the questions @BlueMiku

OpenStudy (bluemiku):

i know i see it now :)

OpenStudy (anonymous):

yeah i have a hard time with story problems so i need help on them

OpenStudy (bluemiku):

ok i will try my best on helping you :)

OpenStudy (only4u):

first one is easy dear

OpenStudy (anonymous):

okay thankyou :)

OpenStudy (bluemiku):

(84 X2/7 = 12 X2 = 24 gallons) the first one

OpenStudy (only4u):

2/7 tank=84*2/7

OpenStudy (anonymous):

okay yeah i guess that makes sense thanks George cut 2/3 of the pie and put that giant piece on his plate. Then he ate 1/4 of that piece. what fraction of the original pie did George eat?

OpenStudy (only4u):

1/4*2/3

OpenStudy (bluemiku):

WOW! that was fast XD

OpenStudy (only4u):

2/3 of x=y 1/4 of y =1/4*2/3 of x

OpenStudy (anonymous):

i still don't really get this one

OpenStudy (only4u):

its very easy .. what u dont get in this sweety?

OpenStudy (anonymous):

like i don't see the steps so i don't really know how you got your answer or what it is

OpenStudy (only4u):

suppose u have 100 chocolates .. u gave me 20% of them .. how much u have left?

OpenStudy (anonymous):

i really don't know

OpenStudy (only4u):

do u know what is the meaning of 20 percent?

OpenStudy (anonymous):

not really

OpenStudy (only4u):

ok.. divide ur pie into 100 hundred equal parts .. give me 20 parts of them .. now the pie i have is the 20% of the total pie

OpenStudy (anonymous):

its almost 1am and i need to get some sleep can we finish it tomorrow

OpenStudy (only4u):

ok

OpenStudy (anonymous):

goodnight

OpenStudy (only4u):

gn